# CLIMB-3D: Continual Learning for Imbalanced 3D Instance Segmentation

### v1

@@ 4.3 Designing Incremental Scenarios
Frequency Scenarios (Split_A). This scenario acknowledges that datasets are often labeled based on the frequency of category occurrences. To accommodate this, we propose a split where the model initially learns from the most frequent categories and subsequently incorporates the less frequent ones in later stages. By prioritizing the training of frequently occurring categories, the model can establish a strong foundation before expanding its knowledge to handle rarer categories.
2
Semantic Scenarios (Split_B). In real-world environments, objects may exhibit similarities in appearance, and then moved to different environments, the model may encounter new objects that does not share similar semantic characteristics with previously seen categories. To address this, we introduce the Split_B  scenario. It involves grouping categories based on their semantic labels and incrementally training the model to handle these groups. This allows the model to generalize its knowledge across semantically similar categories, facilitating adaptation to new objects with similar characteristics. Unlike the Split_A  scenario, this scenario may include both frequent and infrequent categories within the same task.
3
@@
Random Scenarios (Split_C). In some cases, data labeling is based on the availability of objects rather than specific criteria. To account for this scenario, we design the Split_C  scenario. This scenario represents a completely random setting where each task can have any class, leading to varying degrees of class imbalance. By exposing the model to such diverse and imbalanced scenarios, we aim to enhance its ability to handle real-world situations where the availability of labeled data is unpredictable.
@@
By designing these three incremental learning scenarios, we aim to provide a more realistic representation of object distributions, frequencies, and dynamics encountered in the real world.
@@ 5 Experiments
We evaluate our method across three scenarios specifically designed to capture the complexities of real-world settings, where new categories emerge incrementally alongside class imbalance, as discussed in Section 4.3. Experimental results demonstrate that our approach effectively handles the gradual introduction of new classes and mitigates the impact of class imbalance within these scenarios. The following sections detail the datasets, evaluation metrics, incremental scenarios, and implementation procedures, followed by a thorough analysis and comparison of the results.
5.1 Experimental Setup
Datasets. We evaluate our method using the ScanNet200 dataset [rozenberszki2022language], which includes 200 object categories and exhibits inherent class imbalance, making it well-suited for simulating and evaluating real-world scenarios. Additionally, we benchmark our method against existing incremental learning approaches using the original ScanNet dataset [dai2017scannet] in a semantic segmentation setting. We follow the standard training and validation splits as defined in prior works.
Evaluation Metrics. We evaluate our method using mean Average Precision (mAP), a standard metric for 3D instance segmentation that provides a comprehensive measure of segmentation quality, accounting for both precision and recall. For comparison with existing semantic incremental learning approaches, we report mean Intersection over Union (mIoU), which measures the overlap between predicted and ground truth instances, offering a detailed evaluation of segmentation accuracy. To assess the model’s ability to mitigate catastrophic forgetting in continual learning scenarios, we use the Forgetting Percentage Points (FPP), as defined in [liu2023continual]. This metric quantifies performance degradation by measuring the accuracy drop between the initial and final training phases, on the categories observed in the first training phase.
Incremental Scenarios. As discussed in Section 4.3, we design three incremental scenarios: Split_A, Split_B, and Split_C, each consisting of three tasks, which are grouped based on object occurrence frequency, semantic similarity, and random grouping, respectively. In Split_A, the frequency of object categories progressively decreases through the tasks. This scenario follows the head, common, and tail splits present in the ScanNet200 dataset, with class distributions of 66-68-66 in each split. In Split_B, we partition the classes into 74-50-76 based on semantic similarity, which is calculated using the CLIP [radford2021learning] text encoder, followed by clustering using K-Means. Finally, in Split_C, the classes are shuffled and split into three sets, resulting in 67-67-66 categories per split. These scenarios allow for a comprehensive evaluation of our approach under varying conditions, facilitating a deeper understanding of its performance and generalization in diverse real-world settings.
Implementation Details. We utilize the transformer-based model for 3D instance segmentation proposed in [Schult23ICRA], designed to iteratively attend to hierarchical feature representations. The model processes the 4 coarsest levels of a ResNet-based U-Net backbone across three iterations, progressively refining from coarse to fine, resulting in $L=12$ transformer decoder layers. Each transformer decoder layer shares weights across iterations and consists of a standard transformer layer utilizing self-attention and masked cross-attention mechanisms. The feature backbone employed is Minkowski Res16UNet34C [choy20194d].
Training Details. We adopt the data augmentation, hyperparameters, and training strategy described in [Schult23ICRA]. For joint training (Row 1, Table 4), the model is trained for 600 epochs using the AdamW optimizer [loshchilov2017decoupled] with a one-cycle learning rate scheduler [smith2018superconvergence], and results are evaluated on the entire validation set. In incremental training, we retain the same hyperparameters, adjusting only the number of epochs, and we use a memory buffer size of 50 scenes. Training is divided into three phases, introducing one split per phase across the three designed scenarios. After each phase, the resulting model is evaluated on all classes encountered up to that point.
5.2 Results and Discussion
Table 1: Comparison between the baseline and proposed method with mAP25, mAP50, and mAP, which is after training for all the stages. We also report the FPP metric.
@@ 5.2 Results and Discussion
Split_C CLIMB-3D 31.59 26.78 18.93 9.10 7.89
To evaluate our proposed method, we conduct a comparative analysis using exemplar replay (ER) for instance segmentation and [Yang_2023_CVPR] for semantic segmentation as baselines. As shown in Table 1, our method, which integrates exemplar replay, knowledge distillation, and an imbalance correction (IC) module, achieves notable improvements over the baseline in terms of mAP and FPP. Specifically, in the Split_A  scenario, our approach significantly enhances overall performance. We observe an improvement of 19.23%, 31.05%, and 12.28% for $\text{mAP}_{50}$, $\text{mAP}_{25}$, and overall mAP, respectively, while reducing forgetting by 47.86% as measured by $\text{mAP}_{50}$. For the Split_B  scenario as well, CLIMB-3D demonstrates a consistent performance boost over the baseline, significantly improving mAP and reducing forgetting, which is lowered to 5.52% compared to 46.21% in the baseline for AP50. Likewise, in the Split_C  scenario, CLIMB-3D enhances both learning efficiency and forgetting reduction, achieving a performance of 26.78% in $\text{mAP}_{50}$  and reducing forgetting by 20.95% compared to baseline. These results across scenarios underscore the effectiveness of our approach.
Table 2: Comparison with previous method on semantic segmentation on ScanNet V2 dataset. We report the mIoU metric for the evaluated models.
@@
Yang et al. [Yang_2023_CVPR] 34.16 13.43 28.98
@@
Although our method focuses on segmenting individual objects (instance segmentation), we also demonstrate its performance in semantic segmentation by presenting a comparative analysis with existing methods for class-incremental semantic segmentation on the ScanNet V2 dataset (Table 2). Using our predicted labels, we assign each point the label corresponding to the highest confidence mask and exclude background labels (floor and wall), as these are not part of the object-level segmentation. Following the dataset splits established by [Yang_2023_CVPR], we report results for both training phases. Our proposed method achieves a substantial improvement over prior methods, with a gain of 35.23% in Phase 1 and approximately 19.1% in Phase 2. Overall, our method reaches a mIoU of 59.38%, significantly outperforming previous baselines, which achieve a lower mIoU of around 30%.
@@
We extend the analysis from Table 1 to Table 3 to highlight the impact of our proposed method on individual splits across various scenarios. The results clearly demonstrate that our model consistently retains knowledge of previous tasks better than the baseline. For Split_A , our model shows improvement throughout the phase. In Phase 3 of (s2), although both the baseline and our method exhibit a performance drop, our method reduces forgetting significantly compared to the baseline. The Split_B  scenario, while more complex than Split_A, achieves comparable results due to semantic similarity among classes within the same task. In Phase 2, our model achieves overall all 43.13% $\text{mAP}_{50}$ compared to 24.53% on baseline, a similar trend is observed in Phase 3, where our method not only consistently improves learning but also enhances retention of previous information. After all three tasks, our method achieves an overall performance of 31.56% AP50, compared to 15.07% for the baseline. In the Split_C  scenario, the first-stage model struggles due to the increased complexity introduced by random grouping. In Phase 2, while the baseline focuses on learning the current task, it suffers from severe forgetting of prior knowledge. Conversely, our method balances new task learning with the retention of earlier information. By Phase 3, the model effectively consolidates s1 and maintains strong performance across all task splits. Overall, our proposed method improves mAP by 5.6%.
5.3 Ablation
To assess the effectiveness of each component in our proposed framework, we perform an ablation study. Initially, we establish an upper-bound performance by jointly training the model on the complete dataset using a transformer-based architecture, such as Mask3D [Schult23ICRA], referred to as the Oracle. For the incremental learning setup, we generate training splits according to the scenarios outlined earlier. In this study, we first train the model naively across phases and then sequentially integrate each module to evaluate its individual contribution to performance. Table 4 summarizes the results for the Split_A  scenario, using both the $\text{mAP}_{50}$  and FPP metrics.
Table 4: Ablation study results illustrating the impact of exemplar replay, knowledge distillation, and imbalance correction modules in a three-phase training setup. Each split, representing the subset of data introduced at each phase (p), is labeled as ‘s’ followed by the phase number. The final column, ‘All’, in each phase reports performance across all classes encountered up to that phase. Joint training results (Oracle) are highlighted in gray, while results with all modules combined are marked in orange. The best-performing results are shown in bold.
Row Modules p=1 $\uparrow$ p=2 $\uparrow$ p=3 $\uparrow$ FPP $\downarrow$
s1 s1 s2 All s1 s2 s3 All
1. Oracle - - - - 55.14 30.77 25.30 37.68 -
2. Naïve 56.82 0.00 28.09 14.15 0.00 0.00 19.67 5.80 56.82
3. + ER 56.82 18.51 32.81 25.72 10.38 9.43 24.27 14.28 46.44
4. + KD 56.82 50.00 34.39 42.13 49.78 11.41 26.47 29.28 7.04
5. + IC 56.82 54.67 33.75 44.13 54.19 12.02 26.55 31.05 2.63
Naïve Training. In the naive incremental training setup, where no dedicated modules are incorporated, the model learns the current task but suffers from catastrophic forgetting of the previously learned tasks, as expected. This behavior is evident in row 2, where, upon transitioning to phase 2, the model entirely forgets the classes learned during phase 1. A similar trend is observed in phase 3, and this pattern is also reflected in the FPP metric.
Effect of Exemplar Replay. To mitigate catastrophic forgetting, we incorporate exemplar replay, which stores and replays examples from previous tasks. As shown in row 3, exemplar replay improves average precision by 18.5% for s1 in phase 2 and 10.38% in phase 3. It also reduces forgetting for s2 by 9.43% in phase 3, while slightly improving learning on the current task. However, substantial forgetting persists, as reflected in the FPP metric, highlighting the limitations of exemplar replay alone.
Effect of Knowledge Distillation. The addition of knowledge distillation (KD), which retains a copy of the model from previous tasks, facilitates the preservation of past task knowledge while enabling forward knowledge transfer. As shown in row 4, KD considerably reduces forgetting and boosts performance on the current task. Specifically, for s1, KD improves $\text{mAP}_{50}$ by 31.49% in phase 2 and by 39.40% in phase 3, compared to exemplar replay. Overall, KD leads to a 15.00% increase in performance while reducing forgetting by 39.40% after all tasks have been learned.
Effect of Imbalance Correction. The imbalance correction module addresses the class imbalance in the dataset by re-weighting the teacher model’s predictions during KD based on class frequency. As highlighted in row 5 of Table 4, this addition further improves performance. Specifically, for s1, imbalance correction reduces forgetting by 4.67% and 4.41% in phases 2 and 3, respectively, compared to the results without this module (row 4). For s2, while a slight decrease in current task performance is observed in phase 2, this is likely due to the module’s prioritization of mitigating forgetting less frequent classes in previous tasks. In phase 3, performance on s2 and s3 improves. Overall, imbalance correction significantly reduces forgetting, achieving improvements of 4.41% and 43.81% over KD and exemplar replay, respectively.
6 Conclusion
We address the challenge of class-incremental 3D instance segmentation with class imbalance. We propose an innovative approach that integrates a memory-efficient exemplar replay buffer, knowledge distillation, and a novel imbalance correction module. This framework mitigates the forgetting of rare classes during incremental learning by accounting for the frequency of object occurrences. To enable comprehensive evaluation, we design three incremental learning scenarios, each comprising three phases that reflect real-world dynamics. Our experimental results demonstrate that the proposed framework significantly enhances the learning of new classes while reducing forgetting of previously learned ones. The carefully designed scenarios and framework not only offer a strong baseline but also provide a clear benchmark for future research, laying a foundation for more advanced techniques in class-incremental learning.
Supplementary Material
In this supplementary material, we first demonstrate the performance gains on rare classes achieved by incorporating the IC module in Appendix A. Next, we provide detailed split information for all scenarios, based on class names, in Appendix B. Finally, we present a qualitative comparison between the baseline method and our proposed approach in Appendix C.
Appendix Appendix A Evaluation on Rare Categories
The proposed imbalance correction (IC) module, as detailed in Section 4.2, is designed to address the performance gap for rare classes. To assess its impact, we compare its performance with the framework which has exemplar replay (ER) and knowledge distillation (KD). Specifically, we focus on its ability to improve performance for rare classes, which the model encounters infrequently compared to more common classes.
@@ 6 Conclusion
The results, shown in Table 5 and Table 6, correspond to evaluations on Split_A  for Phase 2 and Phase 3, respectively. In Phase 2, we evaluate classes seen 1–20 times per epoch, while Phase 3 targets even less frequent classes, with observations limited to 1–10 times per epoch.
As illustrated in Table 5, the IC module substantially improves performance on rare classes in terms of $\text{mAP}_{50}$  in Phase 2 of Split_A. For instance, classes like recycling bin and trash bin, seen only 3 and 7 times, respectively, shows significant improvement when the IC module is applied. Overall, the IC module provides an average boost of 8.32%, highlighting its effectiveness in mitigating class imbalance.
@@
Similarly, Table 6 presents results for Phase 3, demonstrating significant gains for infrequent classes. For example, even though the classes such as piano, bucket, and laundry basket are observed only once, IC module improves the performance by 52.30%, 10.40%, and 13.60%, respectively. The ER+KD module does not focus on rare classes like shower and toaster which results in low performance, but the IC module compensates for this imbalance by focusing on underrepresented categories. On average, the addition of the proposed IC module into the framework outperforms ER+KD by 12.13%.
@@
Appendix Appendix B Incremental Scenarios Phases
Table 7 presents the task splits for each proposed scenario introduced in Section 4.3 using the ScanNet200 dataset. The three scenarios, Split_A, Split_B, and Split_C, are each divided into three tasks: Task 1, Task 2, and Task 3. Notably, the order of classes in these tasks is random.
Appendix Appendix C Qualitative Results

### v2

@@ 3.3 Benchmarking Incremental Scenarios
Split_A: Frequency Scenarios. This scenario acknowledges that datasets are often labelled based on the frequency of category occurrences. To accommodate this, we propose a split where the model initially learns from the most frequent categories and subsequently incorporates the less frequent ones in later stages. By prioritizing the training of frequently occurring categories, the model can establish a strong foundation before expanding its knowledge to handle rare categories.
Split_B: Semantic Scenarios. Beyond frequency, semantic similarity is crucial in real-world deployments. While objects often share visual or functional traits, models may encounter semantically different categories in new environment. To simulate this challenge, we introduce the Split_Bscenario. Here, categories are grouped based on their semantic relationships, and the model is incrementally trained on one semantic group at a time. This setup encourages the model to generalise across semantically similar categories and adapt more effectively when exposed to new ones. Unlike the Split_Ascenario, which organises learning based on category frequency, the Split_Bscenario may contain both frequent and infrequent categories within a single task, focusing instead on semantic continuity and transfer.
@@
Split_C: Random Scenarios. In some cases, data labelling is driven by the availability of objects rather than predefined criteria. To capture this, we introduce the Split_Cscenario, which represents a fully random setting where any class can appear in any task, resulting in varying levels of class imbalance. By exposing the model to such diverse and unpredictable distributions, we aim to improve its robustness in real-world situations where labelled data availability is inconsistent.
@@
By designing these incremental scenarios, we aim to provide a more realistic representation of object distributions, frequencies, and dynamics encountered in the real world.
@@ 4 Experiments
Table 1: Comparison between the proposed method and the baseline in the 3DIS setting, evaluated using mAP25, mAP50, mAP, and FPP after training across all phase.
@@
Split_C CLIMB-3D (Ours) 31.59 26.78 18.93 9.10 7.89
Table 2: Comparison between the proposed method and existing baselines in the 3DSS setting on ScanNetV2 [dai2017scannet], evaluated using mIoU.
@@
Yang Yang_2023_CVPR 34.16 13.43 28.98
@@
Table 3: Ablation study showing the impact of each component ER, PLG, and CBR in a three-phase setup. Each split (s) and number corresponds to data introduced at that phase. The All column reports performance over all classes seen so far. Joint training (Oracle) is highlighted in gray, and the full model in orange. Best results are in bold.
Row Modules p=1 $\uparrow$ p=2 $\uparrow$ p=3 $\uparrow$ FPP $\downarrow$
s1 s1 s2 All s1 s2 s3 All
1. Oracle - - - - 55.14 30.77 25.30 37.68 -
2. Naïve 56.82 0.00 28.09 14.15 0.00 0.00 19.67 5.80 56.82
3. + ER 56.82 18.51 32.81 25.72 10.38 9.43 24.27 14.28 46.44
4. + PLG 56.82 50.00 34.39 42.13 49.78 11.41 26.47 29.28 7.04
5. + CBR 56.82 54.67 33.75 44.13 54.19 12.02 26.55 31.05 2.63
4.1 Setup
Datasets. We evaluate our method on the ScanNet200 [rozenberszki2022language], which comprises 200 object categories and exhibits significant class imbalance – making it ideal for simulating and assessing real-world scenarios. In addition, we benchmark our approach against existing incremental learning methods using the ScanNetV2 [dai2017scannet] in a 3DSS setting. For this evaluation, we follow the standard training and validation splits defined in prior works [Yang_2023_CVPR].
Evaluation Metrics. We evaluate our method using mean Average Precision (mAP), a standard metric for 3DIS that provides a comprehensive measure of segmentation quality by accounting for both precision and recall. For comparison with existing 3DSS methods, we report the mean Intersection over Union (mIoU), which quantifies the overlap between predicted and ground truth segments. To assess the model’s ability to mitigate catastrophic forgetting in incremental settings, we use the Forgetting Percentage Points (FPP) metric, as defined in [liu2023continual]. FPP captures performance degradation by measuring the accuracy drop on the initially seen categories between the first and final training phases.
Incremental Scenarios. As discussed in Section 3.3, we design three incremental scenarios – Split_A, Split_B, and Split_C– each with three tasks, based on object frequency, semantic similarity, and random grouping, respectively. For Split_A, we follow the class split provided by rozenberszki2022language, with a class distribution of 66-68-66 across the three tasks. In Split_B, we partition the classes into 74-50-76 based on semantic similarity, which is calculated using the CLIP [radford2021learning] text encoder, followed by clustering using K-Means. Finally, in Split_C, classes are randomly shuffled and split into three sets, with 67-67-66 categories per split. These scenarios comprehensively evaluate our approach’s performance and generalization in varied real-world conditions.
Implementation Details. For all the experiments we use Mask3D [Schult23ICRA] with the Minkowski Res16UNet34C [choy20194d] as a feature backbone. We adopt the data augmentation, hyperparameters, and training strategy described in [Schult23ICRA], with using the AdamW optimizer [loshchilov2017decoupled] and a one-cycle learning rate scheduler [smith2018superconvergence], and results are evaluated on the entire validation set. In incremental training, learning is divided into three phases, introducing one split per phase across the three scenarios described in Sec. 3.3. For each phase, we train for 500, 300, and 300 epochs, respectively, except for the Oracle setting (Row 1 in Tab. 3), which is trained once for 600 epochs. We retain the same hyperparameters throughout, modifying only the number of training epochs. The exemplar replay size is fixed at $|\mathcal{E}^{t}|=50$ scenes.
4.2 Results and Discussion
To evaluate our proposed approach, we construct a baseline using ER [buzzega2020dark] for the 3DIS setting, as no existing incremental baselines are available for this task. In contrast, for the 3DSS setting, where prior incremental baselines do exist, we compare our method against three existing approaches [serra2018overcoming, Yang_2023_CVPR, li2017learning].
Results on CI-3DIS. Tab. 1 compares CLIMB-3D against the ER baseline across three CI-3DIS scenarios, evaluated after all phases using $\text{mAP}_{25}$, $\text{mAP}_{50}$, overall mAP, and FPP. In the Split_Ascenario, characterized by significant distribution shifts, it achieves gains of +19.23%, +16.76%, and +12.28% in $\text{mAP}_{25}$, $\text{mAP}_{50}$, and overall mAP, while drastically reducing forgetting with FPP scores of 3.44% and 2.63%, over 47 points lower than the baseline. Under the Split_Bscenario, with semantically related new categories, improvements of +18.26%, +16.49%, and +12.76% are observed alongside a reduction in forgetting to 8.00% and 5.51% from baseline levels above 42%. For the Split_Cscenario, featuring increasing geometric complexity, it maintains solid gains of +5.94%, +5.70%, and +4.08% across mAP metrics, with forgetting reduced by over 22 points. Overall, these results demonstrate robust forward transfer, minimal forgetting, and stable learning, validating the method’s effectiveness for scalable continual 3D semantic understanding.
Results on CI-3DSS. Although our primary focus is on CI-3DIS, we additionally evaluate our method under the CI-3DSS setting to enable comparison with existing approaches. For this, we adapt our predictions by assigning each point the label corresponding to the highest-confidence mask and exclude background classes (floor and wall), as these are not part of the semantic segmentation targets. Tab. 2 presents a comparison between our method and existing baselines on the ScanNet V2 dataset [dai2017scannet], evaluated using mIoU across two training phases and overall. Despite being designed for instance segmentation, our method generalises effectively to semantic segmentation, achieving substantial gains: +35.23% mIoU in Phase 1 and +19.1% in Phase 2. Overall, it achieves 59.38% mIoU – significantly higher than the $\sim$30mIoU of prior methods – highlighting its robustness and transferability.
For a detailed per-phase analysis, evaluation of rare classes, and qualitative result comparisons, please refer to Appendix Appendix A, Appendix Appendix B, and Appendix Appendix D, respectively, in the supplementary material.
4.3 Ablation
We conduct an ablation study to evaluate the contribution of each component in our framework. As an upper bound, we report the performance of an Oracle model trained jointly on the full dataset. For the incremental setting, we follow the previously defined splits and first train the model naively across phases. We then incrementally add each module to isolate its impact. Tab. 3 presents results on the Split_Ascenario using both $\text{mAP}_{50}$  and FPP metrics.
Naïve Training. When trained naïvely without any dedicated modules, the model suffers from severe catastrophic forgetting, as evident in row 2. The model completely forgets previously learned classes upon entering new phases, with performance dropping to zero for earlier splits and a high FPP of 56.82.
Effect of ER. Adding exemplar replay (row 3) partially alleviates forgetting by maintaining a buffer of past examples. This yields notable gains for s1 in phase 2 (+18.51%) and phase 3 (+10.38%), and also improves s2 in phase 3 (+9.43%). However, the overall forgetting remains substantial (FPP: 46.44), showing ER alone is insufficient.
Effect of PLG. The addition of the pseudo-label generator (PLG), which generates labels for previous classes by retaining a copy of the model from earlier phase, facilitates knowledge retention and forward transfer. As shown in row 4, PLG significantly reduces forgetting and enhances performance on current tasks. For s1, it improves $\text{mAP}_{50}$ by 31.49% in phase 2 and 39.40% in phase 3 over exemplar replay. Overall, PLG yields a 15.00% increase in performance and reduces forgetting by 39.40%.
Effect of CBR. Finally, class-balanced re-weighting (CBR) mitigates class imbalance both during pseudo-labelling and current task learning by adjusting each class’s contribution based on its frequency. As shown in row 5, CBR enhances s1 retention over PLG in phases 2 and 3 (+4.67% and +4.41%) and further improves performance on s2 and s3. It also achieves the lowest FPP of 2.63, reflecting effective forgetting mitigation. Overall, CBR offers the best trade-off between retaining past knowledge and learning new tasks, outperforming PLG and ER by 4.41% and 43.81%, respectively.
5 Conclusion
We address the critical challenge of catastrophic forgetting in class-incremental 3D instance segmentation by proposing a modular framework that combines exemplar replay (ER), a pseudo-label generator (PLG), and class-balanced re-weighting (CBR). Our method incrementally adapts to new classes while preserving knowledge from previous tasks without requiring access to the full dataset. Through extensive experiments on a three-phase benchmark, we demonstrate the individual and combined effectiveness of each component. ER provides a easy knowledge storage, PLG enhances knowledge retention via pseudo-label supervision, and CBR further refines learning by mitigating class imbalance. Together, these modules significantly reduce forgetting and improve overall segmentation performance across all phases. By balancing stability and plasticity, our method contributes to continual 3D scene understanding and establishing a new baseline for future research.
Supplementary Material
@@ 5 Conclusion
Figure 3: Tasks are grouped into incremental scenarios based on object frequency, semantic similarity, and random assignment.  ,  , and   denote different tasks; shapes indicate object categories;   marks the background. Left: Grouped by category frequency. Middle: Grouped by semantic similarity (e.g., similar shapes). Right: Randomly grouped, mixing semantic and frequency variations.
@@
In this supplementary material, we first present the per-phase performance of the proposed CLIMB-3D model in Appendix Appendix A, followed by the analysis of gains on rare classes achieved through the incorporation of the CBR module in Appendix Appendix B. Next, we provide detailed split information for all scenarios, based on class names, in Appendix Appendix C. Finally, we present a qualitative comparison between the baseline method and our proposed approach in Appendix Appendix D.
Appendix Appendix A Per-Phase Performance Analysis
We extend the analysis presented in Tab. 1 (Sec. 4.2 of the main paper) to Tab. 4 to examine the impact of our proposed method across individual splits and phases in different scenarios. The results consistently show that our method outperforms the baseline by better retaining knowledge from earlier tasks.
Split_A. In the Split_Ascenario, our model shows notable improvements across all phases. Particularly in Phase 3, although both methods show a drop in performance for s2, our approach significantly reduces forgetting, maintaining a higher $\text{mAP}_{50}$  across all splits.
Split_B. The Split_Bscenario, while more challenging due to increased semantic overlap, demonstrates our model’s strength in generalisation. In Phase 2, it achieves an overall $\text{mAP}_{50}$  of 43.13% compared to the baseline’s 24.53%. This trend continues in Phase 3, with our model achieving 31.56%, more than double the baseline’s 15.07%, indicating both improved learning and retention.
Split_C. For the Split_Cscenario, where splits are randomly grouped, Phase 1 performance is initially lower. However, in subsequent phases, our method balances stability and plasticity better than the baseline. By Phase 3, it consolidates performance across all splits, improving the overall $\text{mAP}_{50}$to 26.78%, a 5.7% gain over the baseline.
These results demonstrate the robustness of our method in handling continual learning under varied and challenging split strategies.
Appendix Appendix B Evaluation on Rare Categories
The proposed Class-Balanced Re-weighting (CBR) module, as detailed in 3.2, is designed to address the performance gap for rare classes. To assess its impact, we compare its performance with the framework which has exemplar replay (ER) and Pseudo-Label Generator (PLG). Specifically, we focus on its ability to improve performance for rare classes, which the model encounters infrequently compared to more common classes.
The results, shown in Tab. 5 and Table 6, correspond to evaluations on Split_A  for Phase 2 and Phase 3, respectively. In Phase 2, we evaluate classes seen 1–20 times per epoch, while Phase 3 targets even less frequent classes, with observations limited to 1–10 times per epoch.
As illustrated in Table 5, the CBR module substantially improves performance on rare classes in terms of $\text{mAP}_{50}$  in Phase 2 of Split_A. For instance, classes like recycling bin and trash bin, seen only 3 and 7 times, respectively, show significant improvement when the CBR module is applied. Overall, the CBR module provides an average boost of 8.32%, highlighting its effectiveness in mitigating class imbalance.
Similarly, Table 6 presents results for Phase 3, demonstrating significant gains for infrequent classes. For example, even though classes such as piano, bucket, and laundry basket are observed only once, the CBR module improves the performance by 52.30%, 10.40%, and 13.60%, respectively. The ER+PLG module does not focus on rare classes like shower and toaster, which results in low performance, but the CBR module compensates for this imbalance by focusing on underrepresented categories. On average, the addition of the proposed CBR module into the framework outperforms ER+PLG by 12.13%.
Appendix Appendix C Incremental Scenarios Phases
Table 7 presents the task splits for each proposed scenario introduced in Sec. 3.3, and Fig. 3 illustrates them visually for the ScanNet200 [rozenberszki2022language] dataset. The three scenarios – Split_A, Split_B, and Split_C– are each divided into three tasks: Task1, Task2, and Task3. Notably, the order of classes within these tasks is random.
Appendix Appendix D Qualitative Results